ellen wishes to mix candy worth $1.80 per pound with candy worth $2.40 per pound to form 48 pounds of a mixture worth $2.00 per pound. How many pounds of the more expensive candy should she use.
Answer by mananth(16949)

candy worth $1.80 per pound --- let the quantity be x lbs
candy worth $2.40 per pound ---- this will be 48-x lbs
Mix 48 pounds of a mixture worth $2.00 per pound.
...
1.80x+2.4(48-x)=48*2
1.80x+115.2-2.4x=96
0.6x=115.2-96
0.6x=19.2
x=32 lbs of $1.80
...
m.ananth@hotmail.ca